OpenStudy (anonymous):

Please help? 3 1/6 * 2 1/4 Multiply and put answer in simplest form.

OpenStudy (anonymous):

\[3\frac{ 1 }{ 6 } \times 2\frac{ 1 }{ 4 }\]

OpenStudy (anonymous):

5 5/12 im pretty sure

OpenStudy (anonymous):

Convert mixed numbers to improper fractions, multiply, then divide. Correct?

OpenStudy (anonymous):

i just made them have the same denominators ( 2/12 , 3/12 ) and multiplied them (3+2=5) and you would multiply the 3 and 2 which would be 10 not 5

OpenStudy (anonymous):

so 10 5/12

OpenStudy (anonymous):

See I got \[7 \frac{ 3 }{ 24 }\] . Change the mixed numbers to improper fractions, so 3 1/6 to 19/6, and 2 1/4 to 9/4. Multiply the two, and you have 171/24. Divide. 24 goes into 171 seven times with three left over, therefore you are left with the mixed number 7 3/24. My question is multiple choice, and neither of our answers are one of the choices. A) \[7 \frac{ 1 }{ 8}\] B) \[7\frac{ 1 }{ 4 }\] C) \[7\frac{ 1 }{ 3 }\] D) \[7\frac{ 1 }{ 2 }\]

OpenStudy (anonymous):

reduce 3/24 and thats 1/8 so you got it right. its A

OpenStudy (anonymous):

Oh duh

OpenStudy (anonymous):

-face palms-

OpenStudy (anonymous):

its okay i so the same thing lol